OpenStudy (anonymous):

Which function's graph has a period of 2? (answer choices posted inside!)

OpenStudy (anonymous):

answer choices A,B,C,D from top to bottom ***im not too sure, but i think it's either B or C... what do y'all think? :)

terenzreignz (terenzreignz):

If it's sin, cos, sec, csc, then, given that the thing inside the function is \[\Large trig\left(bx+c\right)\] Its period is \[\huge \frac{2\pi}{b}\] Note: You can replace trig by any of the following : sin, cos, sec, csc.

OpenStudy (anonymous):

okay, so sin(bx+c) cos(bx+c) sec(bx+c) csc(bx+c) ??

terenzreignz (terenzreignz):

Yeah, so, you see that b? (b is whatever coefficient x has) 2pi/b gives you the period.

OpenStudy (anonymous):

and with what u put above, i'm thinking the answer is C. y= 3cos pi x since you have the pi in it... and the pi is after the 3cos part... is that right?? :) is it C? :D because of that?

terenzreignz (terenzreignz):

C is y = -4sin 2x :P

OpenStudy (anonymous):

ohh sorry!! lol i meant B.. is the answer B then?? :)

terenzreignz (terenzreignz):

Yeah, it's B, since (T is period) \[\huge T=\frac{2\pi}{\pi}=2\]

OpenStudy (anonymous):

aweosme! :D yayy :D thanks!!!